OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y?

pooja195 (pooja195):

First turn it into slope intercept for y=mx+b

OpenStudy (alexandervonhumboldt2):

lets convert to slope intercept form

OpenStudy (alexandervonhumboldt2):

first subtarct 4x from both sides

OpenStudy (alexandervonhumboldt2):

12-4x=4x-4x-6y can you simplify this?

OpenStudy (anonymous):

yea one sec

OpenStudy (anonymous):

12-4x= -6y

pooja195 (pooja195):

The 4x's cancel out 4x-12=6y then just divide everything by 6 4x/6-12/6=6y/6 y=____x-____

OpenStudy (alexandervonhumboldt2):

yes now divide both sides by -6. -6y/(-6)=(12-4x)/(-6)

OpenStudy (anonymous):

y=2-2/3

pooja195 (pooja195):

indicate the slope

pooja195 (pooja195):

y=2/3x-2 y=mx+b m=slope

OpenStudy (anonymous):

i kinda uck at math so can you simplify i for me

pooja195 (pooja195):

You only need to pick out the number in front of the x :P

pooja195 (pooja195):

y=\(\huge\rm\color{red}{2/3x}\)-2

OpenStudy (alexandervonhumboldt2):

y=mx+b form where m is slope

OpenStudy (anonymous):

so my answer is...

pooja195 (pooja195):

I made it bold in red. You should be able to figure it out.

pooja195 (pooja195):

y=\(\huge\rm\color{red}{2/3x}\)-2

OpenStudy (anonymous):

im sorry for any troubles

pooja195 (pooja195):

You're fine lol :)

OpenStudy (anonymous):

but Idk

OpenStudy (anonymous):

lol

pooja195 (pooja195):

What numbers are in red?

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO

OpenStudy (anonymous):

my answer is 2/3x

pooja195 (pooja195):

YEEESSSS WELL DONE ::DDDDD The slope is just 2/3 though :P

OpenStudy (anonymous):

THANK YOU SO MUCH!

pooja195 (pooja195):

No problem!! :D

OpenStudy (anonymous):

I might ask you for some more thanks

pooja195 (pooja195):

You can just tag me i'll be available to help hopefully : )

OpenStudy (anonymous):

alright, thanks